# Regret Analysis for Randomized Gaussian Process Upper Confidence Bound

### v1

@@ Appendix C Proofs for High Probability Regret Bounds of IRGP-UCB
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a positive semidefinite kernel and $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, and ${\cal X}$ be finite. Pick $\delta\in(0,1)$. Assume that $\zeta_{t}$ follows a shifted exponential distribution with $s_{t}=2\log(|{\cal X}|t^{2}\pi^{2}/(12\delta))$ and $\lambda=1/2$ for any $t\geq 1$. Then, by running IRGP-UCB with $\zeta_{t}$, the cumulative regret can be bounded with probability at least $1-\delta$ as follows:
$R_{T}$ $\leq 2\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}}+2\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}\right)}$
$=O\bigl{(}\sqrt{T\gamma_{T}\log(|{\cal X}|T/\delta)}\bigr{)}.$
@@
From the assumption that $s_{t}=2\log(|{\cal X}|t^{2}\pi^{2}/(12\delta))$ and $\lambda=1/2$, we obtain that $\mathbb{E}[U_{t}]=\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}}$ regarding $U_{t}$ in Lemma 4.3. Furthermore, from Eq. (4.3) in Laurent and Massart (2000), we can obtain
$\Pr\left(\sum_{t=1}^{T}2\log\frac{1}{U_{t}}\geq 2T+2\sqrt{2T\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}}+2\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}\right)\leq\frac{6\delta}{\pi^{2}(T+1)^{2}}.$
Note that $2\log\frac{1}{U_{t}}=s_{t}+\log\frac{1}{\tilde{U}_{t}}$, where $\tilde{U}_{t}\sim{\rm Uni}(0,1)$, follows the exponential distribution with $\lambda=1/2$ and $\sum_{t=1}^{T}2\log\frac{1}{\tilde{U}_{t}}$ follows the chi-square distribution whose degree of freedom is $2T$. Therefore, using union bound, the following events hold with probability at least $1-\delta$,
  1. 1.
    $\forall t\geq 1$ and $\forall\boldsymbol{x}\in{\cal X}$, $|f(\boldsymbol{x})-\mu_{t-1}(\boldsymbol{x}_{t})|\leq\zeta_{t}^{1/2}\sigma_{t-1}(\boldsymbol{x}_{t})$,
  2. 2.
    $\sum_{t=1}^{T}\zeta_{t}\leq 2Ts_{T}+2T+2\sqrt{2T\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}}+2\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}$.
Assuming the above events hold, we can obtain the regret upper bound as follows:
@@
$\leq 2\sum_{t=1}^{T}\zeta_{t}^{1/2}\sigma_{t-1}(\boldsymbol{x}_{t})$
$\leq 2\sqrt{\sum_{t=1}^{T}\zeta_{t}\sum_{t=1}^{T}\sigma_{t-1}^{2}(\boldsymbol{x}_{t})}$
$\leq 2\sqrt{C_{1}\gamma_{T}\sum_{t=1}^{T}\zeta_{t}}$
@@
$\leq 2\sqrt{C_{1}\gamma_{T}\left(2Ts_{T}+2T+2\sqrt{2T\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}}+2\log\bigl{(}(T+1)^{2}\pi^{2}/(6\delta)\bigr{)}\right)}.$
For the inequalities, we used the assumed event 1, Cauchy–Schwartz inequality, Lemma 5.4 in Srinivas et al. (2010), and the assumed event 2 in order. ∎
Appendix D Auxiliary Lemmas
@@
Let ${\cal X}\subset[0,r]^{d}$ be a compact and convex set, where $r>0$. Assume that the kernel $k$ satisfies the following condition on the derivatives of sample path $f$. There exists the constants $a,b>0$ such that,
@@
Then, from Assumption 2.1, we obtain several lemmas. First, we show the upper bound of the supremum of the partial derivatives. This result is tighter than Lemma 12 in (Kandasamy et al., 2018). Note that, in the proof of Lemma 12 in (Kandasamy et al., 2018), there is a typo that $\Pr(L\geq t)$ is bounded above by $a\exp\{t^{2}/b^{2}\}$, which should be $ad\exp\{-t^{2}/b^{2}\}$.
Lemma D.1.
@@
From the assumption, using union bound, we obtain the bound of the probability,
@@
Lemma D.2.
Let $f\sim{\cal G}{\cal P}(0,k)$ and Assumption 2.1 holds. Let ${\cal X}_{t}\subset{\cal X}$ be a finite set with each dimension equally divided into $\tau_{t}=bdru_{t}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}$ for any $t\geq 1$. Then, we can bound the expectation of differences,
@@
Lemma D.3 (in Lemma 5.2 of (Srinivas et al., 2010)).
@@
Proof.
Let $R\sim{\cal N}(0,1)$. Then, we obtain the following:
$\Pr(R>c)$ $=\int_{c}^{\infty}\frac{1}{\sqrt{2\pi}}\exp(-r^{2}/2){\rm d}r$ (40)
$=\exp(-c^{2}/2)\int_{c}^{\infty}\frac{1}{\sqrt{2\pi}}\exp(-r^{2}/2+c^{2}/2){\rm d}r$ (41)
$=\exp(-c^{2}/2)\int_{c}^{\infty}\frac{1}{\sqrt{2\pi}}\exp(-(r-c)^{2}/2-rc+c^{2}){\rm d}r$ (42)
$=\exp(-c^{2}/2)\int_{c}^{\infty}\frac{1}{\sqrt{2\pi}}\exp(-(r-c)^{2}/2-c(r-c)){\rm d}r$ (43)
[…]
$=\frac{1}{2}\exp(-c^{2}/2).$ (45)

### v3

@@ Appendix C Proofs for High-Probability Regret Bounds of IRGP-UCB
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a positive semidefinite kernel and $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, and ${\cal X}$ be finite. Pick $\delta\in(0,1)$. Assume that $\zeta_{t}=2\log(1/U_{t})$, where $\{U_{t}\}_{t\geq 1}$ is a sequence of mutually independent random variables that satisfy $\mathbb{E}[U_{t}]=\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}}$ and $\Pr(U_{t}\in(0,1))=1$ for all $t\geq 1$. Then, by running the randomized GP-UCB with $\zeta_{t}$, the cumulative regret can be bounded with probability at least $1-\delta$ as follows:
$\Pr\left(\forall T\geq 1,R_{T}\leq 2\sqrt{C_{1}\gamma_{T}\sum_{t=1}^{T}\zeta_{t}}\right)\geq 1-\delta,$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
@@
We can transform the regret as follows:
@@
[…]
Next, we show the following corollary:
Corollary 4.1.
Let $f\sim{\cal G}{\cal P}(0,k)$, where $k$ is a positive semidefinite kernel and $k(\boldsymbol{x},\boldsymbol{x})\leq 1$, and ${\cal X}$ be finite. Pick $\delta\in(0,1)$. Assume that $\zeta_{t}$ follows a shifted exponential distribution with $s_{t}=2\log(|{\cal X}|t^{2}\pi^{2}/(6\delta))$ and $\lambda=1/2$ for any $t\geq 1$. Then, by running IRGP-UCB with $\zeta_{t}$, the cumulative regret can be bounded with probability at least $1-\delta$ as follows:
$\Pr\left(\forall T\geq 1,R_{T}\leq 2\sqrt{C_{1}\gamma_{T}\left(Ts_{T}+T+2\sqrt{T\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)}+2\log\left(\frac{\pi^{2}T^{2}}{3\delta}\right)\right)}\right)\geq 1-\delta,$
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$.
Proof.
From the assumption that $s_{t}=2\log(|{\cal X}|t^{2}\pi^{2}/(6\delta))$ and $\lambda=1/2$, we obtain that $U_{t}\sim{\rm Uni}(0,\frac{6\delta}{|{\cal X}|t^{2}\pi^{2}})$ and $\mathbb{E}[U_{t}]=\frac{3\delta}{|{\cal X}|t^{2}\pi^{2}}$, where $\zeta_{t}=2\log(1/U_{t})$ as in Lemma 4.3. Moreover, from Theorem 4.5, we obtain
@@
$\Pr\left(\forall T\geq 1,R_{T}\leq 2\sqrt{C_{1}\gamma_{T}\sum_{t=1}^{T}\zeta_{t}}\right)\geq 1-\frac{\delta}{2}.$
Furthermore, since $\sum_{t=1}^{T}\{\zeta_{t}-s_{t}\}$ follows a chi-square distribution with $T$ degrees of freedom, using the union bound and Lemma E.4, we can obtain
$\Pr\left(\forall T\geq 1,\sum_{t=1}^{T}\zeta_{t}\leq Ts_{T}+T+2\sqrt{T\log\frac{\pi^{2}T^{2}}{3\delta}}+2\log\frac{\pi^{2}T^{2}}{3\delta}\right)\geq 1-\frac{\delta}{2}.$
Therefore, taking the union bound concludes the proof. ∎
Appendix D Proof for Expected Regret Bound of GP-UCB with Constant Confidence Parameter
First, we derive the following lemma to show Theorem 4.6:
Lemma D.1.
Let $\{\epsilon_{i}\}_{i\geq 1}$ be a mutually independent sequence of standard normal random variables, that is, $\epsilon_{i}\sim{\cal N}(0,1)$ for all $i\geq 1$. Then, there exists a positive constant $C$ such that
$\Pr\left(\forall t\leq T,\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}\geq-1\right)\geq C.$
Proof.
From the properties of the probability and the union bound, we can see that
$\Pr\left(\forall t\leq T,\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}\geq-1\right)$ $=1-\Pr\left(\exists t\leq T,\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}<-1\right)$
$\geq 1-\sum_{t=1}^{T}\Pr\left(\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}<-1\right)$
$=1-\sum_{t=1}^{T}\Pr\left(\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}>1\right),$
where the last equality is obtained from the symmetry of the Gaussian distribution. Since $\epsilon_{i}\sim{\cal N}(0,1)$ for all $i\geq 1$, the average $\sum_{i=1}^{t}\epsilon_{i}/t\sim{\cal N}(0,1/t)$. Therefore, by using Lemma E.3, we obtain that
$\Pr\left(\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}>1\right)\leq\frac{1}{2}\exp\left(-\frac{t}{2}\right).$
Since $\bigl{(}\frac{1}{2}\exp\left(-\frac{t}{2}\right)\bigr{)}_{t\geq 1}$ is a geometric series with common ratio $\exp(-\frac{1}{2})$ and initial term $\frac{1}{2}\exp(-\frac{1}{2})$, its sum of infinite series converges as follows:
$\sum_{t=1}^{T}\Pr\left(\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}>1\right)$ $\leq\sum_{t=1}^{T}\frac{1}{2}\exp\left(-\frac{t}{2}\right)$
$\leq\sum_{t=1}^{\infty}\frac{1}{2}\exp\left(-\frac{t}{2}\right)$
$=\frac{1}{2\exp(1/2)\bigl{(}1-\exp(-1/2)\bigr{)}}\approx 0.771.$
Consequently, we can obtain the following lower bound of the probability:
$\Pr\left(\forall t\leq T,\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}\geq-1\right)$ $\geq 1-\sum_{t=1}^{T}\Pr\left(\frac{\sum_{i=1}^{t}\epsilon_{i}}{t}>1\right)$
$\geq 1-\frac{1}{2\exp(1/2)\bigl{(}1-\exp(-1/2)\bigr{)}}\approx 0.229.$
Then, we show the following theorem:
Theorem 4.6.
Assume that ${\cal X}=\{\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)}\}$, $\epsilon_{t}\sim{\cal N}(0,1)$ for all $t\geq 1$, and
$\left(\begin{array}[]{c}f(\boldsymbol{x}^{(1)})\\ f(\boldsymbol{x}^{(2)})\end{array}\right)\sim{\cal N}\left(\left(\begin{array}[]{c}0\\ 0\end{array}\right),\left(\begin{array}[]{cc}1&\rho\\ \rho&0.99\end{array}\right)\right),$
where $\rho<1$ is a covariance parameter. Then, if GP-UCB with any constant confidence parameter $\beta_{t}^{(1/2)}=c>0$ runs, then BCR grows linearly, that is,
${\rm BCR}_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\right]=\Omega(T).$
Proof.
We consider the following event:
$E_{T}=\left\{f(\boldsymbol{x}^{(1)})\geq\frac{2\max\{1,c\}}{1-\rho}+1,f(\boldsymbol{x}^{(2)})>f(\boldsymbol{x}^{(1)})+1,\text{ and }\sum_{i=1}^{t}\epsilon_{i}/t\geq-1\text{ for all }t\geq 1\right\}.$
From Lemma D.1 and the independentness between $f$ and $(\epsilon_{i})_{i\geq 1}$, the probability of this event $\Pr(E_{T})>0$. Then, the BCR can be bounded from below as follows:
${\rm BCR}_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\right]$ (25)
$=\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid E_{T}\right]\Pr(E_{T})+\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid E_{T}^{\rm c}\right]\Pr(E_{T}^{\rm c})$ (26)
$\geq\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid E_{T}\right]\Pr(E_{T}),$ (27)
where $E_{T}^{\rm c}$ is a complementary event of $E_{T}$.
We show that, when $E_{T}$ holds, $\boldsymbol{x}_{t}=\boldsymbol{x}^{(1)}$ for all $t\leq T$. First, since the variance of $\boldsymbol{x}^{(1)}$ is larger than that of $\boldsymbol{x}^{(2)}$ in the prior distribution, the first chosen input $\boldsymbol{x}_{1}=\boldsymbol{x}^{(1)}$. Then, we show $\boldsymbol{x}_{t+1}=\boldsymbol{x}^{(1)}$ for all $t\geq 1$. Fix $t\geq 1$ and assume that $\boldsymbol{x}_{i}=\boldsymbol{x}^{(1)}$ for all $i\leq t$. The posterior mean can be derived as follows:
$\mu_{t}(\boldsymbol{x}^{(1)})$ $=\boldsymbol{1}_{t}^{\top}\left(\boldsymbol{1}_{t}\boldsymbol{1}_{t}^{\top}+\boldsymbol{I}_{t}\right)^{-1}\boldsymbol{y}_{t},$
$\mu_{t}(\boldsymbol{x}^{(2)})$ $=\rho\boldsymbol{1}_{t}^{\top}\left(\boldsymbol{1}_{t}\boldsymbol{1}_{t}^{\top}+\boldsymbol{I}_{t}\right)^{-1}\boldsymbol{y}_{t},$
where $\boldsymbol{1}_{t}=(1,\dots,1)^{\top}\in\mathbb{R}^{t}$ and $\boldsymbol{I}_{t}\in\mathbb{R}^{t\times t}$ is the identity matrix. From the Woodbury formula, we obtain
$\left(\boldsymbol{1}_{t}\boldsymbol{1}_{t}^{\top}+\boldsymbol{I}_{t}\right)^{-1}=\boldsymbol{I}_{t}-\frac{\boldsymbol{1}_{t}\boldsymbol{1}_{t}^{\top}}{t+1}.$
Therefore, the posterior means can be transformed as
$\mu_{t}(\boldsymbol{x}^{(1)})$ $=\sum_{i=1}^{t}y_{i}-\frac{t}{t+1}\sum_{i=1}^{t}y_{i}=\frac{t}{t+1}\frac{\sum_{i=1}^{t}y_{i}}{t},$
$\mu_{t}(\boldsymbol{x}^{(2)})$ $=\frac{\rho t}{t+1}\frac{\sum_{i=1}^{t}y_{i}}{t}.$
Thus, the difference between the posterior means is bounded from below as
$\mu_{t}(\boldsymbol{x}^{(1)})-\mu_{t}(\boldsymbol{x}^{(2)})$ $=(1-\rho)\frac{t}{t+1}\frac{\sum_{i=1}^{t}y_{i}}{t}$
$\geq\frac{1-\rho}{2}\frac{tf(\boldsymbol{x}^{(1)})+\sum_{i=1}^{t}\epsilon_{i}}{t}$ $(\because t/(t+1)\geq 1/2),$
$\geq\frac{1-\rho}{2}\frac{\frac{2t\max\{1,c\}}{1-\rho}+1-1}{t}$ $(\because E_{T}\text{ holds})$
$=\max\{1,c\}$
$\geq c$
$=\beta_{t}^{(1/2)}$
$\geq\beta_{t}^{(1/2)}\sigma_{t}(\boldsymbol{x}^{(2)})$ $(\because\sigma_{t}(\boldsymbol{x}^{(2)})\leq 1)$
$\geq\beta_{t}^{(1/2)}\sigma_{t}(\boldsymbol{x}^{(2)})-\beta_{t}^{(1/2)}\sigma_{t}(\boldsymbol{x}^{(1)}).$
Hence, we see that $\boldsymbol{x}_{t+1}=\boldsymbol{x}^{(1)}$ since $\mu_{t}(\boldsymbol{x}^{(1)})+\beta_{t}^{(1/2)}\sigma_{t}(\boldsymbol{x}^{(1)})\geq\mu_{t}(\boldsymbol{x}^{(2)})+\beta_{t}^{(1/2)}\sigma_{t}(\boldsymbol{x}^{(2)})$.
Since $E_{T}$ holds, the optimal point $\boldsymbol{x}^{*}=\boldsymbol{x}^{(2)}$. However, as shown above, $\boldsymbol{x}_{t}=\boldsymbol{x}^{(1)}$ for all $t\geq 1$ if $E_{T}$ holds. Therefore, $\mathbb{E}\left[\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})\mid E_{T}\right]\geq T$. Consequently, since $\Pr(E_{T})>0$, we can conclude that
${\rm BCR}_{T}$ $=\Omega(T).$
Appendix E Auxiliary Lemmas
@@
Let ${\cal X}\subset[0,r]^{d}$ be a compact and convex set, where $r>0$. Assume that the kernel $k$ satisfies the following condition on the derivatives of sample path $f$. There exist the constants $a,b>0$ such that,
@@
Then, from Assumption 2.1, we obtain several lemmas. First, we show an upper bound of the supremum of partial derivatives. This result is tighter than Lemma 12 in [Kandasamy et al., 2018]. Note that, in the proof of Lemma 12 in [Kandasamy et al., 2018], there is a typo that $\Pr(L\geq t)$ is bounded above by $a\exp\{t^{2}/b^{2}\}$, which should be $ad\exp\{-t^{2}/b^{2}\}$.
Lemma E.1.
@@
From the assumption, using the union bound, we obtain a bound on the probability,
@@
Lemma E.2.
Let $f\sim{\cal G}{\cal P}(0,k)$ and Assumption 2.1 holds. Let ${\cal X}_{t}=\{\frac{r}{\tau_{t}+1},\dots,\frac{\tau_{t}r}{\tau_{t}+1}\}^{d}\subset{\cal X}$, where each coordinate takes values from a uniform grid of $\tau_{t}=bdru_{t}\bigl{(}\sqrt{\log(ad)}+\sqrt{\pi}/2\bigr{)}$ points. Then, we can bound the expectation of differences,
@@
Lemma E.3 (in Lemma 5.2 of [Srinivas et al., 2010]).
@@
Lemma E.4 (Eq 4.3 of [Laurent and Massart, 2000]).
Pick $\delta\in(0,1)$. Suppose that $X$ follows a chi-square distribution with $D$ degrees of freedom. Then, the following upper bound for the tail probability holds:
$\Pr\left(X\geq D+2\sqrt{Dc}+2c\right)\leq\exp(-c),$
for any $c>0$. Therefore, the following inequality holds with probability at least $1-\delta$:
$X\leq D+2\sqrt{D\log(1/\delta)}+2\log(1/\delta).$